# Towards Quantifying Long-Range Interactions in Graph Machine Learning: a Large Graph Dataset and a Measurement

### v1

@@ B.1 Baseline Configurations and Resource Consumptions
As mentioned in Section 3, we adopt NeighborLoader from PyG to sample the supporting subgraph for each node. This method implements the sampling strategy of GraphSAGE (Hamilton et al., 2017), which recursively selects $[N_{1},N_{2},\dots,N_{H}]$ neighbors from a node’s 1st, 2nd, …, $H$-th hop neighborhood. Given the grid-like structure of our city networks, we empirically found that the average size of a $16$-hop ego-network is typically around $1k$ nodes, which remains manageable for most GPU devices. Consequently, we set each $[N_{1},N_{2},\dots,N_{H}]$ to $1k$, ensuring that the sampled subgraph covers all nodes in the $16$-hop neighborhood. Furthermore, we use a batch_size of $20k$ (i.e., $20k$ seed nodes) for training and testing. The GPU memory cost and training/testing time are reported in the next section.
Figure 8: GPU memory consumptions when setting hops $H$ equals to # layers $L$ at $H\in[2,4,8,16]$.
Resource monitor.
The experiments are conducted on a cluster of 8 NVIDIA L40 GPUs, each with 48 GB memory. For reference, we monitor the peak GPU memory consumption of models under each setting by running only one job at a time, where we set batch size=$20k$ and vary $L=H\in[2,4,8,16]$. The results are shown in Figure 8, in which we can observe that SGFormer, despite having a medium model size among the baselines, occupies the most memory in all cases due to its attention module.
Figure 9: Average training time per epoch with a batch size of $20k$ seed nodes in NeighborLoader.
Meanwhile, we also record the training/testing clock time per epoch in Figure 9 and Figure 10 respectively. The results suggest that SGFormer generally costs a longer computation time than other GNN baselines at both training and testing, but such difference is insignificant on London. We hypothesize that, since London is at least twice larger than the other networks, it requires more iterations to finish one epoch under the same batch_size=$20k$. This implies that the NeighborLoader needs to sample more ego-networks during training/testing, which is time-consuming when the underlying graph is large.
Figure 10: Average testing time per epoch with a batch size of $20k$ seed nodes in NeighborLoader.
B.2 Experimental Limitations
Neighborhood sampling.
@@ B.2 Experimental Limitations
As shown in Figure 4, when fixing the number of layers at $L=16$, the trend from $H=8$ to $H=16$ remains relatively flat, indicating an incremental performance gain. We attribute this phenomenon to the overlapped subgraphs from batched sampling method NeighborLoader, where the ego-networks of different seed nodes often share common supporting nodes. Consequently, if the number of layers $L$ exceeds $H$, the model can capture information from more distant hops beyond $H$, rather than being limited to a maximum of $H$ as expected. This effect is further amplified when using a large batch size (i.e., a large number of seed nodes) or a large number of hops (i.e. a large neighborhood size), as it increases the chance of overlapping ego-networks. Since the node label approximates eccentricity based on its 16-hop neighborhood, there is minimal additional information beyond the 16th hop. As a result, the performance remains similar between these two settings.
Hyper-parameter selection.
We acknowledge that the results reported in Table 2 may not reflect the models’ optimal performance, as we only considered models with $L=H=\{2,4,8,16\}$ and a fixed hidden_size of 32 in our experiments. It is likely that the baselines could achieve better performance with tuned hyperparameters (e.g., learning rate, weight decay, dropout, hidden size, etc.). Nevertheless, we emphasize once again that the primary goal of this work is to analyze the performance trends of each model when varying the number of layers and the number of hops in the sampled neighborhood. This approach incurs a high computational cost on large graphs, which restricts us from conducting extensive hyperparameter optimization, and we leave this potential exploration for future work.
Appendix C Computational Complexity
@@
where $\rho(v,u)$ is the length of shortest path from node $v$ to node $u$. Given the large scale of our city networks, we employ a stochastic approximation that samples $10k$ nodes to further reduce the computational cost, which can be finished in under 1 hour for all models and networks.
Appendix D Proofs
D.1 Proofs for Theoretical Results in Section 5
@@
Proof of Proposition 14.
@@
Furthermore, as $\mathcal{G}$ is connected $\forall x:d^{\mathcal{G}}_{x}\geq 1$ so $\sum d^{\mathcal{G^{\prime}}}_{x}>n(1+\gamma)$. Therefore:
@@
Proof for Theorem 15.
Given Proposition 14 and the correspondence between the eigenvalues of $\bm{S}_{adj}$ and $\bm{L}^{\mathcal{G}}_{sym}$, the result follows immediately from (Chung, 1997, Lemma 1.14). ∎
D.2 Proofs for Theoretical Results in Section 6
@@
Observe that: $p_{i}(1-p_{i})>0$ for $10<p_{i}<$ 1 and $\partialderivative{p_{i}}{z_{i}}\to 0$ as $p_{i}\to 0$ or $p_{i}\to 1$. For $j\neq i$:
@@
Proof of Proposition 6.1.
Let $f:\mathbb{R}^{D}\to\mathbb{R}$ be a differentiable function with $f^{\prime}(\bm{X})\neq 0$ at some $\bm{X}$. For simplicity set $D=1$ and let $\bm{X}=x$ (scalar case). Define a model by
$k(x)=k_{1}(x)+k_{2}(x),$ (46)
$k_{1}(x)=f(x),$ (47)
$k_{2}=-f(x).$ (48)
We first observe that for any $x$,
$k(x)=f(x)-f(x)=0.$ (49)
By differentiability, the derivative is
$k^{\prime}(x)=k_{1}^{\prime}(x)+k_{2}^{\prime}(x)=f^{\prime}(x)-f^{\prime}(x)=0.$ (50)
Thus, a perturbation in $x$ has no net effect (i.e., zero influence) on $k(x)$. However, we note that there are two branches: $k_{1}^{\prime}(x)$ with derivative $f^{\prime}(x)$, $k_{2}^{\prime}(x)$ with derivative $-f^{\prime}(x).$ If we assess sensitivity by taking the absolute value of each derivative (i.e., the magnitude of each influence), we obtain
$\left|f^{\prime}(x)\right|\quad\text{and}\quad\left|-f^{\prime}(x)\right|=\left|f^{\prime}(x)\right|,$ (51)
so that the total influence would be
$\left|k_{1}^{\prime}(x)\right|+\left|k_{2}^{\prime}(x)\right|=2\left|f^{\prime}(x)\right|.$ (52)
Definition D.1 (Smooth Hypersurface in $\mathbb{R}^{D}$).
A hypersurface in $\mathbb{R}^{D}$ is a subset defined locally as the zero set of a continuously differentiable function $f:\mathbb{R}^{D}\to\mathbb{R}$ such that the gradient $\nabla f$ is nonzero at almost every point where $f=0$. That is, if we have an equation of the form $f(\bm{\theta})=0$, where $\bm{\theta}$ is a vector of parameters, and if $\nabla f(\bm{\theta})\neq 0$ generically, then $f=0$ defines a (locally) $(D-1)$-dimensional manifold, which is a hypersurface.
Proof of Theorem 6.3.
For exact cancellation to hold assuming $\bigoplus$=$\sum$ = ∑, the following sum must be identically zero while the individual terms remain non-zero. Since $\psi$ and $\phi$ are differentiable and parameterized by $\bm{\theta}$, each Jacobian term is a smooth function of $\bm{\theta}$. The equation:
@@
defines a level set of smooth functions, a hypersurface (or a set of lower-dimensional submanifolds) in parameter space. The solution set of a nontrivial smooth function has Lebesgue measure zero unless it is identically zero across all $\bm{\theta}$, which is not the case here. Furthermore, since $\bm{\theta}$ is drawn from an absolutely continuous distribution (such as Gaussian or uniform), the probability of exactly selecting a parameter that lies on this hypersurface is zero at initialization. Thus, exact cancellation of Jacobian contributions is an event of measure zero in the space of our idealized single-layer MPNN: this can naturally be extended to multiple layers. Note that in this proof we have assumed the Jacobian sum is not identically zero by construction, unlike in Proposition 6.1 where the function was explicitly constructed to ensure cancellation. ∎

### v2

@@ C.3 Ablation Studies
Concretely, our strategy is implemented with NeighborLoader from PyG, which recursively selects $N_{1},N_{2},\dots,N_{H}]$ neighbors from a node’s 1st, 2nd, …, $H$-th hop neighborhood. Given the grid-like structure of our city networks, the size of the neighborhood does not explode with $H$, and we empirically found that the average size of a $16$-hop ego-network is typically around $1k$ nodes, which remains manageable for most GPU devices. Therefore, we sample all nodes inside the $H$-hop neighborhood, and use a batch_size of $20k$ (i.e., $20k$ seed nodes) for training and testing.
The results are presented in Figure 10, where we can observe that every baseline, at a fixed model architecture, shows a consistent upward trend on all four city networks. These results indicate that long-range dependency, rather than the parameter size, is the primary factor that contributes to the improvement of the model’s performance.
@@
One may notice the trend from $H=8$ to $H=16$ remains relatively flat. We attribute this phenomenon to the overlapped subgraphs sampled from NeighborLoader, where the ego-networks of different seed nodes often share common supporting nodes. Consequently, if the number of layers $L$ exceeds $H$, the model can capture information from more distant hops beyond $H$, rather than being limited to a maximum of $H$ as expected. This effect is further amplified when using a large batch size (i.e., a large number of seed nodes) or a large number of hops (i.e. a large neighborhood size), as it increases the chance of overlapping ego-networks. As a result, the performance remains similar between these two settings. Note that this phenomenon is not a contradiction to our main claims, but rather a limitation, as we can not create perfect 8-hop subgraphs.
Figure 9: Ablation results under hidden_size=[16, 32, 64, 128] and num_layers=[2, 4, 8, 16] on Paris, Shanghai, and Los Angeles. The patterns are all consistent with our findings in Section 3.
C.4 Summary of Baseline Results
For reference, we summarize the best baseline results on our City-Networks in Table 5. Note that due to the global attention module in GraphGPS, it is highly memory-consuming and fails to fit into our GPU memory (4848GB), even with hidden_size=$16$ under num_layers=$16$.
Table 5: Baseline results on City-Networks.
Baseline Type Paris Shanghai Los Angeles London
MPNNs MLP $25.5\pm 0.4$ $28.4\pm 0.6$ $24.1\pm 0.5$ $27.9\pm 0.1$
ChebNet $54.1\pm 0.2$ $66.5\pm 0.1$ $61.4\pm 0.4$ $54.7\pm 0.2$
GCN $53.2\pm 0.3$ $62.1\pm 0.2$ $58.3\pm 0.3$ $50.1\pm 0.7$
GraphSAGE $54.6\pm 0.2$ $68.3\pm 0.5$ $61.4\pm 0.3$ $55.4\pm 0.2$
GAT $51.1\pm 0.3$ $68.0\pm 0.5$ $59.5\pm 0.3$ $52.0\pm 0.3$
GCNII $51.3\pm 0.2$ $61.5\pm 0.4$ $56.0\pm 0.3$ $48.2\pm 0.3$
DropEdge $48.2\pm 0.2$ $60.8\pm 0.4$ $55.5\pm 0.3$ $45.0\pm 0.3$
GTs GraphGPS $52.1\pm 0.6$ $63.0\pm 0.5$ $59.8\pm 0.5$ OOM
Exphormer $55.1\pm 0.8$ $70.2\pm 0.4$ $63.8\pm 0.6$ $49.5\pm 0.4$
SGFormer $52.0\pm 0.8$ $64.1\pm 0.3$ $60.1\pm 0.7$ $48.3\pm 0.3$
Appendix D Computational Complexity in R
@@
where $d(v,u)$ is the length of shortest path from node $v$ to node $u$. Given the large scale of our city networks, we employ a stochastic approximation that samples $10k$ nodes to further reduce the computational cost. For reference, the calculation finishes under 30 minutes for all baselines on all four City-Networks using a single NVIDIA RTX 3090 GPU with 48 AMD Ryzen 3960XCPU cores.
Appendix E Limitations in Experiments
Total Influence on large dense graphs.
During our experiment, we found that on large dense graphs, especially the one with small diameters such as ogbn-arxiv ($|V|=169k$, $diam=25$), it is difficult to compute our Total Influence measurement, as $\bar{N}_{H}$ in Equation (13) quickly converges to $N$ as $H$ increases. In this case, if the model also happens to have complex architectures (e.g., GraphGPS), the computation of $R$ and $T_{h}$ will therefore become impractical. We believe this constitutes a limitation of our proposed measurement, where future works may explore more effective methods for dense graphs based on stochastic sampling.
Appendix F Proofs and Justifications in Section 5
F.1 Details of Linearization
A Graph Convolutional Network (GCN) layer with input features $\bm{X}^{(l)}$, learnable weights $\bm{W}^{(l)}$ for layer $l$, and non-linear pointwise activation function $\sigma$ is defined as:
$\bm{\tilde{S}}_{adj}=\bm{\tilde{U}}\bm{\Lambda}\bm{\tilde{U}}^{T},$ (16)
$(\bm{\tilde{S}}_{adj})^{k}=\bm{\tilde{U}}\bm{\Lambda}^{k}\bm{\tilde{U}}^{T},$ (17)
F.2 Proofs for Theoretical Results in Section 5
Lemma F.1 (Eigenvalue complementarity of normalized operators).
$\lambda_{N+1-i}(\bm{{S}}_{adj})=1-\lambda_{i}(\bm{L}_{sym})$ (20)
@@ F.2 Proofs for Theoretical Results in Section 5
This complementarity implies that when the normalized algebraic connectivity $\lambda_{2}(\bm{L}_{sym})$ is small, the second largest positive eigenvalue of $\bm{{S}}_{adj}$ must be close to 1. However, in our context, we are interested in the algebraic connectivity of the normalized graph Laplacian that would correspond to the normalized adjacency operator, $\bm{\tilde{S}}_{adj}$, introduced in Section 5.1, instead of that of $\bm{{S}}_{adj}$, motivating the following:
Proposition F.2 (Self-loops decrease algebraic connectivity of the original graph, from Section 5.2).
Assume a connected graph $\mathcal{G}$ with more than two nodes. For all $\gamma>0$,
Proof of Proposition 4 (same as Proposition 22).
@@
Furthermore, as $\mathcal{G}$ is connected, each node has degree of at least 1 – $\forall x:d^{\mathcal{G}}_{x}\geq 1$. As the graph has more than two nodes, one node must have degree of at least 2. So $\sum d^{\mathcal{G^{\prime}}}_{x}>n(1+\gamma)$, and therefore:
@@
Proof of Theorem 5.
Given Lemma 20 (correspondence between the eigenvalues of $\bm{S}_{adj}$ and $\bm{L}^{\mathcal{G}}_{sym}$) and Proposition 4, seeing that $diam(\mathcal{G})\geq 4$ means that the graph has more than two nodes, the result follows immediately from (Chung, 1997, Lemma 1.14). ∎
F.3 Rationale in Theorem 5
The second largest eigenvalue of $\bm{\tilde{S}}_{adj}$ in terms of magnitude is either $\lambda_{N-1}$ or $\lambda_{1}$ . We explicitly consider the case when it is $\lambda_{N-1}$. As $\lambda_{N-1}$ approaches 1, the rate of convergence to the limiting state $\bm{\tilde{u}}_{N}\bm{\tilde{u}}_{N}^{T}$ decreases exponentially with the number of layers. For any layer $l$, the difference from the limiting state can be expressed as:
[…]
where $||\cdot||$ denotes the spectral norm and the last equality follows from that the spectral norm of a diagonal matrix being equal to the largest absolute value of its diagonal entries. Thus, when $\lambda_{N-1}$ is close to 1, more layers are required to achieve the same level of convergence to the limiting state. Since $\lambda_{N-1}<1$, taking powers of $\lambda_{N-1}$ will eventually converge to zero, but the rate of this convergence slows dramatically as $\lambda_{N-1}$ approaches 1. Hence, because graphs with a large diameter and low maximum degree have a lower algebraic connectivity due to reduced inter-component connectivity, GNNs operating on such graphs will be less susceptible to over-smoothing when processing node features.
Nevertheless, our analysis demonstrates that large diameters and sparse connectivity generally reduce the likelihood of over-smoothing. This insight motivates our proposal of benchmark datasets with significantly larger diameters than those currently used in the literature (see Table 1). The underlying hypothesis is that reduced likelihood of over-smoothing in high-diameter, sparse networks enables the possibility of GNNs learning representations that capture long-range dependencies when necessary.
Appendix G Jacobian-based Influence Score Motivation
Here, we provide additional motivation behind the proposed measurement for quantifying long-rangeness in Section 4.
G.1 Intuition behind Jacobian-based Influence Score
where $\bm{X}$ is the original (unperturbed) input feature matrix, and $\delta\bm{e}_{uj}$ is an infinitesimal perturbation in the $j^{th}$ component of the feature vector at node $u$ (a standard basis vector in the node feature space). This means the more positive the Jacobian entry is, the more a positive perturbation of the features at node $u$ and component $j$ will increase the logits at node $v$ and component $i$ at the final layer. In other words, the Jacobian entry being positive or negative means that the logits are pushed up or down. Given that, after applying the softmax function, the probabilities at a given data point increase monotonically with the logits at that point, we can consider both positive and negative influences as actual influence and only focus on the absolute value (i.e., sensitivity rather than direction).
@@ G.1 Intuition behind Jacobian-based Influence Score
Observe that: $p_{i}(1-p_{i})>0$ for $0<p_{i}<1$ and $\partialderivative{p_{i}}{z_{i}}\to 0$ as $p_{i}\to 0$ or $p_{i}\to 1$. For $j\neq i$:
@@
The above is a well-known fact and we do not claim novelty.
G.2 Potential Limitations: Output Cancellation
Proposition G.1 (Absolute Jacobian sensitivity may over-estimate influence).
There exists a model $h_{v}$, where the combined effect of changes in input variables on the output is zero (i.e., the sum of the partial derivatives is zero), while the sum of the absolute values of the individual partial derivatives is nonzero.
Proof of Proposition G.1.
Consider a graph with three nodes $u,v,w$, where $v$ is the focal node for our calculation. Let input and output features be scalars $\mathbf{X}_{u}=x_{u}$ on $\mathbf{H}_{v}=h_{v}$ (nodewise binary classification where using a single logit as input to a sigmoid function is possible). The influence is $I(v,u)=\left|\frac{\partial h_{v}}{\partial x_{u}}\right|$, and similarly for $\mathbf{X}_{w}=x_{w}$, $I(v,w)=\left|\frac{\partial h_{v}}{\partial x_{w}}\right|$. Assume the model learns the function $h_{v}=x_{u}-x_{w}$ and that $x_{u}\approx x_{w}$, then following the definition in the main text:
$T_{h}(v)=I(v,u)+I(v,w)=\left|\frac{\partial h_{v}}{\partial x_{u}}\right|+\left|\frac{\partial h_{v}}{\partial x_{w}}\right|=|1|+|-1|=2,$ (44)
while $h_{v}=0.$ In fact,
$\frac{\partial h_{v}}{\partial x_{u}}+\frac{\partial h_{v}}{\partial x_{w}}=1+(-1)=0$ (45)
holds for all $(x_{u},x_{w})$; no small-difference assumption on $x_{u}$ and $x_{w}$ is required.
This demonstrates that the sum of the absolute values of the Jacobian entries can be non-zero, even when the net effect on the output is zero. The key is the opposing nature of their influence, not their specific values. The example sets $x_{u}$ and $x_{w}$ as approximately equal to highlight that the net output can be small (or changes to it can be small) while the individual influences are significant. ∎
Although this is a valid concern, we next show that for Message Passing Neural Networks (MPNNs), at least at initialization, such cancellation does not happen.
Definition G.2 (Message-Passing Neural Network layer).
Definition G.3 (Smooth Hypersurface in $\mathbb{R}^{D}$).
A hypersurface in $\mathbb{R}^{D}$ is a subset defined locally as the zero set of a continuously differentiable function $f:\mathbb{R}^{D}\to\mathbb{R}$ such that the gradient $\nabla f$ is nonzero at almost every point where $f=0$. That is, if we have an equation of the form $f(\bm{\theta})=0$, where $\bm{\theta}$ is a vector of parameters, and if $\nabla f(\bm{\theta})\neq 0$ generically, then $f=0$ defines a (locally) (D1)(D-1)-dimensional manifold, which is a hypersurface.
Theorem G.4 (Measure-zero of exact cancellation at MPNN initialization).
Proof of Theorem G.4.
For exact cancellation to hold assuming $\bigoplus=\sum$, the following sum must be identically zero while the individual terms remain non-zero. Since $\psi$ and $\phi$ are differentiable and parameterized by $\bm{\theta}$, each Jacobian term is a smooth function of $\bm{\theta}$. The equation:
@@ G.2 Potential Limitations: Output Cancellation
defines a level set of smooth functions, a hypersurface (or a set of lower-dimensional submanifolds) in parameter space. The solution set of a nontrivial smooth function has Lebesgue measure zero unless it is identically zero across all $\bm{\theta}$, which is not the case here. Furthermore, since $\bm{\theta}$ is drawn from an absolutely continuous distribution (such as Gaussian or uniform), the probability of exactly selecting a parameter that lies on this hypersurface is zero at initialization. Thus, exact cancellation of Jacobian contributions is an event of measure zero in the space of our idealized single-layer MPNN: this can naturally be extended to multiple layers. Note that in this proof we have assumed the Jacobian sum is not identically zero by construction, unlike in Proposition G.1 where the function was explicitly constructed to ensure cancellation. ∎
Lastly, it is worth noting an alternative perspective on interpreting influence measures derived from the sum of absolute Jacobian entries, particularly when considering potential cancellation effects as demonstrated in Proposition G.1. One can indeed argue that the measure’s utility lies precisely in its capacity to quantify the magnitude of sensitivity to inputs from individual distant nodes or pathways, irrespective of whether these influences ultimately negate one another in contributing to the final output. From this stance, focusing on the sum of absolute values reveals the underlying structure and strength of potential long-range dependencies (the information channels themselves) that might be obscured by observing only the net effect. This interpretation, therefore, hinges on defining dependency or interaction based on the existence and intensity of these information flow pathways, rather than strictly on their final, combined impact on a node’s prediction.
Appendix H Additional Definitions in Section 6
Definition H.1 (Standard lattice in $\mathbb{R}^{D}$).
Definition H.2 ($h$-hop shells).
Definition H.3 (Quasi-isometric graphs).
Appendix I Proofs for Theoretical Results in Section 6